Use the information provided below to construct equations relating consumption and saving to disposable income:

income consumption saving tax
25 million 5 million 15 million 5 million
1
Expert's answer
2015-08-19T00:00:44-0400
The consumption equation is:
C = Ca + c*(Y - T), where C - consumption, c - marginal propensity to consume, Ca - autonomous consumption, Y - income, T - tax.
c = C/Y
C = Ca + 5/25*(25 - 5) = Ca + 0.2*(Y - 5) = Ca - 1 + 0.2Y